*p&p ZP z/sl2z+ ,+ ,x) z og zX-1+1 * w
Answer #1

172 trt12_y2 I (x + y + 22 ²3 d2 de de 9-51-72_y2 et 2= II 1-2²-42 2²22 += x-x2-y2 x²+4²+ z² = 22 Champen spherical Let x² +let net Celo = it - fhodoz at = t do = P. 2 = 27 m [1+] 6X7